# On the uniform convexity of the squared distance

### v1

@@ 4 Shadow sequences
which is notable for having been rediscovered in the 2000s by Terence Tao during his work in ergodic theory [31], and popularized through his blog [30], where he treated the case of the monotone convergence theorem [^2] I said ‘rediscovered’, because even this monotone convergence case in itself had already been studied by Georg Kreisel in the 1950s, as it can be seen in the very paper [16, pp. 49–50] which introduced the program now known as proof mining. and asked for suggestions for a proper name for the property. The name that ultimately stuck was metastability, as suggested by Jennifer Chayes, and one may then speak of a ‘rate of metastability’ as being an upper bound on the $N$ in the metastability statement, given in terms of the $\varepsilon$ and the $g$ (and perhaps some other parameters of the problem at hand). Thus, we will state all our convergence results in terms of rates of metastability.
To express such rates, we introduce some notations. For all $f:\mathbb{N}\to\mathbb{N}$, we define $\widetilde{f}:\mathbb{N}\to\mathbb{N}$, for all $n$, by $\widetilde{f}(n):=n+f(n)$ and $f^{M}:\mathbb{N}\to\mathbb{N}$, for all $n$, by $f^{M}(n):=\max_{i\leq n}f(i)$; in addition, for all $n\in\mathbb{N}$, we denote by $f^{(n)}$ the $n$-fold composition of $f$ with itself.
@@ 4.2 Fejér monotone sequences
The following is the generalization from Hilbert spaces to complete $UCW$-hyperbolic spaces of [4, Proposition 5.7] (see also [28, Lemma 3.2]).
@@
Let $\varepsilon$, $n$, $m$ be as stated. Using the projection property and Theorem 3.1 (since $d(P_{S}x_{m},x_{m})\leq b$ and $d(P_{S}x_{n},x_{m})\leq b$), we have that
@@
We now show that the sequence $(d^{2}(P_{S}x_{n},x_{n}))$ is a nonincreasing sequence in $[0,b^{2}]$. Let $n\in\mathbb{N}$. If $P_{S}x_{n}=P_{S}x_{n+1}$, then, using the Fejér monotonicity, $d(P_{S}x_{n+1},x_{n+1})=d(P_{S}x_{n},x_{n+1})\leq d(P_{S}x_{n},x_{n})$. If $P_{S}x_{n}\neq P_{S}x_{n+1}$, then, by the previous property, we have that
@@ 4.3 Quasi-Fejér monotone sequences
The following is the generalization from complete CAT(0) spaces to complete $UCW$-hyperbolic spaces of [24, Lemma 5].
@@
Let $\varepsilon$, $n$, $m$ be as stated. Using the projection property and Theorem 3.1 (since $d(P_{S}x_{m},x_{m})\leq C$ and $d(P_{S}x_{n},x_{m})\leq C$), we have that
@@
$2\psi_{\eta}(C,\varepsilon)\leq\psi_{\eta}(C,\varepsilon)+C\cdot\frac{\psi_{\eta}(C,\varepsilon)}{4C}=\frac{5}{4}\psi_{\eta}(C,\varepsilon),$

### v2

@@ 4 Shadow sequences
which is notable for having been rediscovered in the 2000s by Terence Tao during his work in ergodic theory [33], and popularized through his blog [32], where he treated the case of the monotone convergence theorem [^2] I said ‘rediscovered’, because even this monotone convergence case in itself had already been studied by Georg Kreisel in the 1950s, as it can be seen in the very paper [18, pp. 49–50] which introduced the program now known as proof mining. and asked for suggestions for a proper name for the property. The name that ultimately stuck was metastability, as suggested by Jennifer Chayes, and one may then speak of a ‘rate of metastability’ as being an upper bound on the $N$ in the metastability statement, given in terms of the $\varepsilon$ and the $g$ (and perhaps some other parameters of the problem at hand). Thus, we will state all our convergence results in terms of rates of metastability.
To express such rates, we introduce some notations. For all $f:\mathbb{N}\to\mathbb{N}$, we define $\widetilde{f}$$\mathbb{N}\to\mathbb{N}$, for all $n$, by $\widetilde{f}(n):=n+f(n)$ and $f^{M}:\mathbb{N}\to\mathbb{N}$, for all $n$, by $f^{M}(n):=\max_{i\leq n}f(i)$; in addition, for all $n\in\mathbb{N}$, we denote by $f^{(n)}$ the $n$-fold composition of $f$ with itself.
@@ 4.2 Fejér monotone sequences
The following is the generalization from Hilbert spaces to complete $UCW$-hyperbolic spaces of [4, Proposition 5.7] (see also [30, Lemma 3.2]).
@@
Let $\varepsilon$, $n$, $m$ be as stated. Using the projection property, the convexity of $S$ and Theorem 3.1 (since $d(P_{S}x_{m},x_{m})\leq b$ and $d(P_{S}x_{n},x_{m})\leq b$), we have that
@@
We now show that the sequence $(d^{2}(P_{S}x_{n},x_{n}))$ is a nonincreasing sequence in $0,b^{2}]$. Let $n\in\mathbb{N}$. If $P_{S}x_{n}=P_{S}x_{n+1}$, then, using the Fejér monotonicity, $d(P_{S}x_{n+1},x_{n+1})=d(P_{S}x_{n},x_{n+1})\leq d(P_{S}x_{n},x_{n})$. If $P_{S}x_{n}\neq P_{S}x_{n+1}$, then, by the previous property, we have that
@@ 4.3 Quasi-Fejér monotone sequences
The following is the generalization from complete CAT(0) spaces to complete $UCW$-hyperbolic spaces of [26, Lemma 5].
@@
Let $\varepsilon$, $n$, $m$ be as stated. Using the projection property, the convexity of $S$ and Theorem 3.1 (since $d(P_{S}x_{m},x_{m})\leq C$ and $d(P_{S}x_{n},x_{m})\leq C$), we have that
@@
$0<2\psi_{\eta}(C,\varepsilon)\leq\psi_{\eta}(C,\varepsilon)+C\cdot\frac{\psi_{\eta}(C,\varepsilon)}{4C}=\frac{5}{4}\psi_{\eta}(C,\varepsilon),$